if an iguana is 4.5 feet long,and his tail is two-thirds his body length, how long is his tail.
Answer by stanbon(75887)

if an iguana is 4.5 feet long,and his tail is two-thirds his body length, how long is his tail.
---
4.5 feet = 48 + 6 inches = 54 inches
----
(2/3)54 = 2*18 = 36 inches = 3 feet (tail length)
====================
Cheers,
Stan H